Habitable Zone Planet Finder

DevelopmentsLinks
Sept-Oct
2015
Painting of the interior and anti-room floor
Nov 30
2017
First Light
July
2018
Instrument shutdown for vaccum and lfc maintenance
Nov
2018
Shared Risk Science begins for the 18-3 period
Dec
2018
Normal Science Operations begins (the 19-1 period)
Dec
2018
New TSL key words added to support setup method and comparison lamp
Dec 25
2018
LFC comb went offline and began using the Etalon
Jan 6
2019
LFC repaired
Oct 23 6
2019
LFC briefly lost phase-lock but still functional
Nov 11-12
2019
LFC maintenance to improve efficiency; also replaced cal shutter controllers and a burned out UNe lamp.
Dec 11-14
2019
LFC offline; restored around 06:00 on 2019-12-14 UT.
Dec 15-19
2019
LFC offline; restored before the night of 2019-12-20 UT.
Jan 21
2020
Fiber agitator motor replaced
Feb 10
2020
HPF offline for 1 night (2020-02-11 UT) due to saturation from IHMP work.
Jul 14
2020
HPF observing low-precision targets only for 1 night (2020-07-15 UT) after LN2 vacuum jacketed lines were pumped.
Aug 5
2020
HPF unavailable for most of the night (2020-08-06 UT) after a software issue resulted in the detector being powered down for an hour or two.
Oct 16
2020
HPF observing low-precision targets only for 2 nights (2020-10-17 & 18 UT) after LN2 vacuum jacketed lines were pumped.
Dec 18
2020
HPF observing low-precision targets only for 1 night (2020-12-19 UT) after LN2 vacuum jacketed lines were pumped.
throughout 2021 Multiple nights of low-precision targets only after pumping VJ lines (further occurrences not noted on this page).
Jun 14
2021
HPF observing low-precision targets only while LFC is having problems, starting 2021-06-15 UT for a few nights.
Aug 2-6
2021
HPF LFC team on site doing maintenance and improvements including replacing UNe and Alpha Bright lamps; no impact to precision or observations
Nov 1-10
2021
HPF offline for science while instrument team investigates problem with increasing bias pedestal level in detector
Dec 17
2021
HPF fiber scrambler motor replaced during the night after it fails while in use
Apr 7
2022
HPF fiber scrambler motor replaced during the afternoon when it stopped turning the belts
May 20 to June 14
2022
HPF offline for engineering work (electronics and dewar work inside the spectrograph enclosure; instrument itself remained sealed)
June 15
2022
HPF resumes low-precision observations as instrument stabilizes
Aug 1
2022
HPF resumes high-precision observations




Last updated: Fri, 29 Jul 2022 03:09:22 -0500 stevenj



Overview

The Telescope

Technical Overview

Object Observability

Performance

Instruments

LRS2

LRS2 Summary

LRS2 Details

LRS2 Fiber Layout and Position Angle

LRS2 Throughput

LRS2 Observing details

VIRUS

VIRUS - Summary

VIRUS - Setting up on targets

VIRUS - Misc details

VIRUS - Throughput and sensitivity

VIRUS - IFU Layout and Target Placement

VIRUS - Tiling observations

Habitable Zone Planet Finder (HPF)

HPF Details

HPF Setting up on targets

HPF Throughput and Exposure Meter

HPF Data Reductions

HRS-2 (in development)

HRS-2 Summary

HRS-2 Details

HRS-2 configurations

HRS-2 exposure meter

HRS-2 Position Angle and Fiber layout

HRS-2 Throughput

Old Instruments

HRS - old

LRS - old

MRS - old

Program Preparation

Web Management System